### v1

# Asymmetric Influence of the Amplitude-Dependent Tune Shift on the Transverse Mode-Coupling Instability

Abstract
At the 3 GeV ring of the MAX IV Laboratory, a fourth generation ring-based synchrotron light source, an asymmetric influence of the sign of the amplitude-dependent tune shift (ADTS) on the transverse mode-coupling instability (TMCI) has been observed. Measurements of the instability, in dedicated single-bunch experiments at low chromaticity, revealed a significant dependence of the dynamics of the instability above threshold on the sign of the ADTS. While for a negative sign of the ADTS the crossing of the instability does not lead to a loss of beam current, a positive sign results in the loss of 40% or more of the beam current at the threshold. In order to investigate the observed asymmetry systematic measurement of beam dynamics above the threshold have been conducted in combination with particle tracking simulations with mbtrack2 and theoretical calculations of the Landau damping including the ADTS. The findings indicated that this effect could become relevant in future low-emittance electron storage rings.
pacs:
no PACS numbers yet
I Introduction
The transverse mode-coupling instability (TMCI) in electron storage rings is a single-bunch transverse instability and an important collective effect which can limit the parameter space for stable operation. Especially in fourth generation light sources, this collective effect can strongly influence the achievable operation parameters. The instability depends on many beam parameters like the natural bunch length, the chromaticity and the tunes. The connection with the amplitude-dependent tune shift (ADTS) was in the past mainly investigated in the interest of mitigation of the instability by Landau damping [1, 2, 3], as the required betatron tune spread can among other sources come from the ADTS.
First studies of the influence of the ADTS on the TMCI for the 3 GeV ring of the MAX IV Laboratory where presented by several of the authors in [4, 5]. While in the case of MAX IV the TMCI does not affect the standard operation, it is nevertheless important to characterize and further investigate such instabilities as with the continuous push towards more extreme operation modes and beam parameters new effects and interaction between parameters can arise and become relevant for the mitigation of such instabilities in future machines.
Dedicated, systematic experiments have now been conducted in single-bunch operation showing an asymmetric influence of the sign of the ADTS on the dynamics of the vertical TMCI above threshold. While the ADTS does not seem to significantly affect the threshold current it changes the behaviour of the bunch above threshold. For values of the ADTS close to zero a partial beam loss is observed when the threshold current is crossed while slowly increasing the bunch current. For a positive sign of the ADTS this partial beam loss occurs up to higher values of the ADTS than for a negative sign, resulting in a significant asymmetry in the encountered beam loss above threshold.
The beam dynamics above the TMCI threshold, can be studied at the ADTS values where not beam loss occurs. This is used to investigate the reason behind the observed asymmetry. It is observed that, as expected, the instability leads to a blow up of the bunch size and strong center-of-mass oscillations. For negative ADTS values, and at high current at high positive ADTS values, these oscillations are additionally amplitude-modulated with a much lower frequency showing a sawtooth shaped pattern.
This paper will compare measurements conducted at the MAX IV Laboratory with dedicated simulations and theoretical considerations. The measurements include systematic scans of the instability threshold and the occurring current-loss as well as time-resolved measurements of the beam dynamics, in this case the center-of-mass motion and the transverse bunch size, above the threshold current.
I.1 Transverse Mode-Coupling Instability
The transverse mode-coupling instability can arise when the current dependent tune shift due to the transverse impedance leads to a coupling of the coherent betatron tune with one of the neighboring head-tail mode frequencies (typically the mode -1 with a separation of $-\nu_{s}$, the synchrotron tune). The TMCI, which occurs at zero chromaticity, has, opposed to the head-tail instability, a well defined threshold current at which the growth rate increases abruptly. Figure. 1 shows the simulated mode coupling at zero chromaticity.
The theory of Landau damping in combination with transverse mode-coupling has been developed by Chin [6]. It is characterized by a dispersion integral written as
[…]
where $V$ is the complex coherent tune of the instability to be found, $J$ is the action of betatron oscillation, $\nu_{\beta}$ and $\nu_{s}$ are the betatron and synchrotron tunes respectively, $f$ is the normalized charge distribution in $J$ of the bunch and $\Delta\nu(J)$ is the amplitude-dependent tune shift (ADTS) giving rise to the tune spread.
For the theoretical calculations a Gaussian charge distribution is assumed:
$f(J)=\frac{1}{2\pi\langle J\rangle}e^{-\frac{J}{\langle J\rangle}}\,,$ (2)
where $\langle J\rangle$ is the average action of the particle ensemble. For our purposes, the tune shift is assumed to be proportional to the action $J$ resulting in the following definition where $b$ is the amplitude-dependent tune shift coefficient:
$\omega\left(J\right)=b\cdot J+\omega\left(0\right)\rightarrow b=\frac{\Delta\omega}{J},$ (3)
In the following, when a value for the ADTS coefficient is given, it refers to $b$ and has the unit $\left[b\right]=\nicefrac{{1}}{{\textrm{m}}}$, if not stated otherwise.
Stability is then determined by evaluating the integral for solutions whose growth rate $\mathrm{Im}(V)$ is small and positive. This allows the standard exponential integral $\mathrm{Ei}$ to be used under said approximation. The dispersion relation then evaluates to
$I_{m}=-\frac{1}{b\langle J\rangle}\left[1-\zeta e^{-\zeta}(\mathrm{Ei}(\zeta)-i\pi)\right]\>\mathrm{where}\>\zeta>0$ (4)
where
$\zeta=\frac{V-\nu_{\beta}-m\nu_{s}}{b\langle J\rangle}\,.$ (5)
To include radiation damping, we could evaluate Eq. 1 for $\mathrm{Im}(V)=+1/\tau_{x}$ where $\tau_{x}$ is the radiation damping time, although the integral would then have to be evaluated numerically, slowing down the calculation significantly. In practice, the radiation damping does not make a large difference in the case of the TMCI so it is neglected.
The inverse of the dispersion relation is subtracted from the diagonal elements of Chin’s scaled coupling matrix $\nu_{s}\mathbf{M}_{nl}^{mk}$ for head-tail and mode-coupling instabilities as given by Eq. 2.44 in [6]. Solutions are determined numerically by equating the determinant of the resulting matrix to zero:
$\mathrm{det}(I^{-1}_{m}\delta_{ml}\delta_{nk}-\nu_{s}\mathbf{M}_{nl}^{mk})=0$ (6)
where $\delta_{ij}$ is the Kronecker delta. In practice, there are two unknowns left to determine: the tune spread at $\langle J\rangle$ ($\Delta\nu(\langle J\rangle)=b\langle J\rangle$) and the coherent frequency of oscillation $\mathrm{Re}(V)$.
Figure 1: Simulated coherent beam spectrum of the bunch at zero chromaticity and zero ADTS, showing the tune shift with increasing bunch current and the resulting mode coupling at the TMCI threshold at 3.5 mA.
II Experimental Setup
Table 1: Beam Parameters during Measurements
Parameter Value
Beam energy / GeV 3.0
Circumference / m 528
RF frequency / MHz 99.931
Harmonic number 176
RF voltage / kV 864
Synchrotron freq. / Hz 830
Synchrotron tune 0.00146
Vertical tune 16.275
Horizontal tune 42.2
The measurements presented in this paper were performed with the machine parameters given in Table 1 if not stated otherwise. Only a single bunch was stored in the machine to be able to use diagnostics that are not bunch-resolved such as beam position monitors (BPM) and the synchrotron light monitor (SLM). To ensure that only a pure single bunch is filled, great care has been taken to clean out residual charge from the other buckets. The total acceleration voltage in the main cavities was set to a fixed value for better comparability of different measurements. Due to the usage of a single bunch and therefore a low absolute beam current, the passive Landau cavities are not elongating the bunch. Likewise, all insertion device gaps were opened for the sake of comparability. The reference orbit for the orbit correction was set to zero-orbit without any beam line bumps laid in, to use an optic close to the design optics used in the simulations. During the measurements itself the orbit correction was only in use while changing to new settings, e.g. changing chromaticity or ADTS and was afterwards switched off. When changing settings, it was also checked that the tunes were at the standard working point. The vertical chromaticity was reduced via sextupole magnets in the vertical plane while the horizontal chromaticity remained at the value for standard operation. For adjustments of the ADTS the three octupole magnet families can be used [7].
Figure 2: BPM signal showing the amplitude modulation caused by the TMCI on the vertical center-of-mass position as function of turns.
II.1 ADTS - measurement and control
The center-of-mass (COM) motion is measured via the turn-by-turn data from the beam position monitors (BPM). The position at each turn can be written out for $2^{16}$ consecutive turns. As the measurements where taken in single-bunch operation, this gives the COM position of the bunch at each turn at every BPM. Additionally, the data observed by the bunch-by-bunch system is saved with a turn-by-turn resolution as well. The coherent tunes can therefore be calculated from the Fourier transform of this data.
The amplitude-dependent tune shift (ADTS) was measured by kicking the bunch in each transverse plane individually with increasing amplitude while detecting the center-of-mass movement on the turn-by-turn BPM data for both transverse planes. The tune was calculated via the Numerical Analysis of Fundamental Frequencies (NAFF) algorithm [8, 9] based on the first 100 turns after each kick. The resulting tunes for the different kick amplitudes show the tune shift as a function of the center-of-mass displacement $\hat{x}$ at each BPM position. For the conversion between the measured maximal amplitude $\hat{x}$ at each BPM and the action $J$, it is assumed that at maximum displacement $\hat{x}$ the action $J$ can be calculated via the corresponding value of the beta-function $\beta_{\textrm{s}}$ at the position of measurement $s$, in this case the position of each BPM:
$J=\frac{\hat{x}^{2}_{\mathrm{s}}}{\beta_{\mathrm{s}}}.$ (7)
Figure. 3 shows the near linear dependence of the tune shift on the action $J$ in an example measurement.
2
Figure 3: Measurement of the betatron tune as function of the action $\langle J\rangle$ in the vertical plane, showing the linear dependence giving the amplitude-dependent tune shift coefficient $b$.
To ease the operation during the experimental scans of the ADTS value, a response matrix $M$ was measured for the resulting change of the ADTS caused by changes to two of the octupole families. This matrix allows fast calculation of the necessary change $\Delta I_{\mathrm{oct,u}}$ in octupole current for a requested change of $\Delta b_{\mathrm{u}}$ in ADTS coefficient, with $u$ representing the planes $x$ and $y$, without the need for intermediate measurements.
$\begin{pmatrix}\Delta I_{\mathrm{oct,x}}\\ \Delta I_{\mathrm{oct,y}}\end{pmatrix}=M^{-1}*\begin{pmatrix}\Delta b_{\mathrm{x}}\\ \Delta b_{\mathrm{y}}\end{pmatrix}$ (8)
Nevertheless, after arriving at a new ADTS value and checking, and if necessary correcting, the chromaticity and the tunes, and before conducting dedicated measurements, the ADTS value was measured to ensure accuracy.
II.2 bunch size
The bunch size is measured at the two diagnostic beam lines [10] via synchrotron light monitors (SLM) with interferometric source point imaging. Synchrotron radiation in the visible wavelength range is detected with CMOS cameras after passing though a double-slit for the horizontal plane and a diffraction obstacle in the vertical plane. The beam sizes can be calculated from the interferometric visibility in the resulting interference pattern. During the instability the bunch size is blown up to such a degree that the interference pattern is not visible anymore (Fig. 4b) and the distribution is fitted by a Gaussian. The required exposure time of the cameras to gather enough intensity, does not allow for turn-by-turn detection. Nevertheless, the exposure time is short enough ($\approx$ 1  ms) to be able to resolve the time structure of the characteristic amplitude modulation observed on the center-of-mass position (see Fig. 2) caused by the dynamics above the instability threshold. In this case, it has to be taken into account that during the exposure time the camera integrates over the observed center-of-mass oscillations providing a superposition of the center-of-mass motion and the interference pattern containing the beam size information. Furthermore, it has do be asserted that no residual bunches are present during these measurements. As visible in Fig. 4, even a small amount of charge in additional residual bunches around the main single bunch, in this case approximately 5% of the charge, can significantly influence the observed spot profile on the SLMs. As these low-charge residual bunches are below the instability threshold and therefore stable, their light shows the classical interference pattern. The intensity of the focused pattern is therefore overshadowing the smeared out spot profile from the unstable main bunch.
Figure 4: Synchrotron light spot of a vertically unstable bunch measured at a diagnostic beam line for interferometric bunch size measurement. a) The additional, residual low-current bunches are stable and show the typical interferometric pattern on top of the unstable main bunch. b) After cleaning the residual bunches, only the vertically unstable single bunch is visible.
II.3 synchronised measurements
The readout of the cameras can be triggered so that synchronised images can be taken with respect to the turn-by-turn center-of-mass motion measured with the BPMs. The synchronization was aligned with triggered kicks to the beam which can be observed in both systems (BPMs and SLMs). The timing between the camera acquisition and BPM acquisition is chosen such that the camera’s exposure time window lies roughly at three-quarters of the BPM measurement window of $2^{16}$ turns ($\approx 115$ ms). By this, the center-of-mass movement is known for some time before and after the bunch size measurement. The alignment accuracy depends on the camera exposure time used and is in the presented measurements better than $1$ ms.
In case of multiple such measurement sets being taken during the instability with the characteristic amplitude modulation on the center-of-mass movement (see Fig. 2), the repetitive behaviour seen on the BPMs can be used to overlay multiple measurement sets aligned by this pattern. This will provide a “sampled” image of the changes in the light spot observed on the SLM cameras. In other words, due to the measurement trigger not being synchronised to the instability dynamics, different phases of the amplitude modulation are sampled with every measurement set taken and the repetitiveness of the amplitude modulation can be used to reconstruct a time resolved image. The spot size is the result of the superposition of the blown up bunch size and the center-of-mass oscillation within the exposure time window.
III Simulation tool
To simulate the beam dynamics observed, especially above the threshold current, particle tracking with the mbtrack2 [11] python code was performed.
A broadband resonator was used for the vertical impedance with a shunt impedance of 200 kOhm/m at the resonant frequency of 11.5 GHz and a quality $Q$ of 1 [12]. The mbtrack2 simulations also included a longitudinal impedance (732 Ohm at 6 GHz with $Q=1$ [13]) to account for bunch lengthening with increasing bunch current. mbtrack2 allows for the optics parameter to be read-in from an AT lattice file using pyAT. The RF voltage was set to the same value as in the measurements (see Table 1). The tune shift contribution by the ADTS is calculated in mbtrack2 based on the action $J$ (see Sec. I.1 Eq. 3). Intra-Beam Scattering (IBS) is not yet implemented in mbtrack2. The shown simulations were conducted with 50000 macro-particles and were run on the COSMOS cluster of LUNARC at Lund University.
Figure 5: Simulated TMCI threshold at a chromaticity of $0.05$, $b=1000$ and different energy spreads of 80%, 90%, 100% and 110% showing the expected increase in threshold for increased energy spread.
Figure 6: left: Single bunch threshold currents during injection shown as a function of ADTS coefficient measured at chromaticities of 0, 0.05 and -0.15. right: Single bunch threshold currents simulated in mbtrack2 (including bunch lengthening by addition longitudinal impedance) for chromaticities of 0, 0.05, -0.15.
IV results
In the following the measurement and simulation results will be presented side by side and grouped by the different beam properties affected by the instability. The measurements were conducted in the vertical plane. Besides the threshold current, the beam loss at threshold, the bunch position and size as well as the betatron tune shift with current below and above the instability threshold are discussed. Additionally, theoretical calculations on the Landau damping in combination with the transverse mode-coupling instability will be discussed in context of the observed asymmetry with respect to the sign of the amplitude-dependent tune shift.
IV.1 Instability Threshold
When studying an instability, the threshold current is a very important parameter as it is the limit up to what current stable operation is possible.
During the experimental investigations of the TMCI it was observed, that the threshold current changes depending on the beam conditions while reaching the threshold. For example, the observed threshold current was lower when the beam current slowly decayed while the beam was unstable, compared to the threshold current observed when charge was injected into a stable beam. Additionally, within a certain bunch current range, it was possible to stabilize an unstable beam with the bunch-by-bunch feedback system and the beam remained stable after switching off the feedback. Furthermore, the instability could be triggered by excitations or kicks to the beam even below the injection threshold, but not below the decaying threshold. In summary, a hysteresis effect was observed for the experimental TMCI threshold current, where a stable beam shows a higher threshold current than an already unstable or excited beam.
A possible explanation for the observed hysteresis in threshold is found when considering the effects of Intra-Beam Scattering (IBS). For beams with small transverse emittances, IBS can lead, amongst other things, to an increase in energy spread. This can be mitigated by increasing the vertical emittance either via coupling or with vertical excitations of the beam. With respect to the TMCI, IBS would have the following effect. For a stable beam the vertical emittance is small and the energy spread is increased by IBS. An increased energy spread results in an increase of the theoretical TMCI threshold, as the current-dependent tune shift is inversely proportional to the bunch length which again is proportional to the energy spread [14]. As soon as the beam becomes unstable, either by crossing the (higher) threshold or by excitation, the vertical emittance increases and the effect of the IBS is reduced leading to a reduction in energy spread. The lower energy spread finally results in a lower TMCI threshold current. This results in a hysteresis of the instability threshold depending on whether the threshold is measured starting with a stable or an unstable beam.
As the mbtrack2 simulations do not include IBS this hysteresis can not be directly simulated. Nevertheless, simulations with the energy spread manually set to different values show the expected dependence of the threshold current on the energy spread (see Fig. 5).
For the following studies the threshold during injection was selected as it can be quickly measured reliably and accurately compared to the other thresholds. Furthermore, the disturbance to the stored beam caused by the used Multipole Injection Kicker (MIK) is known to be very small [15], so the observed threshold during injection should be very close to the theoretical threshold (including IBS) if the charge in a stable beam is slowly increased.
The left hand side of Fig. 6 shows the thresholds measured during injection for different ADTS coefficients. These thresholds where determined by injecting (using the MIK) into a single bunch and observing the center-of-mass movement on the BPMs. As soon as the center-of-mass movement movement grew unstable the injection was stopped and all charge in residual bunches from a non-perfect single bunch injection was cleaned. The resulting threshold currents differ as expected depending on the chromaticity. To separate the TMCI from the head-tail instability [16], the measurements where conducted either at a vertical chromaticity of zero or nearly zero chromaticity (0.05) in contrast to a chromaticity of $\approx 1.1$ during standard operation. The thresholds for both chromaticity values are very similar and lie around 2.8 mA. Additionally, measurements were conducted at a slightly negative vertical chromaticity of $-0.15$. As expected during operation with a positive momentum compaction factor and a negative chromaticity (e.g. [14]), they show a much lower threshold current of around 1 mA. The same is visible in the simulated thresholds shown on the right hand side of Fig. 6. The simulated thresholds for a chromaticity of zero and 0.05 lie both at around 3.45 mA and are higher than seen in the measurements by about 0.5 mA. At the same time, the simulated threshold for the slightly negative chromaticity matches the measurements at around 1 mA.
[…]
IV.2 Beam Losses at Thresholds
Figure 7: Current loss in percent at the TMCI threshold as function of the amplitude-dependent tune shift at chromaticities of 0, 0.05 and -0.15.
A significant influence of the value and sign of the ADTS coefficients can be observed in the amount of charge lost when the instability threshold is crossed during injection. Figure 7 shows the beam loss in percent for different values of the ADTS coefficient with a chromaticity close to zero or with slightly negative chromaticity values. For negative ADTS coefficients up to nearly zero ($\approx-500/\mathrm{m}$ ) no beam loss is encountered at all when crossing the instability threshold during injection. This is already noteworthy as it shows, that the instability is not destructive even though it leads to strong center-of-mass oscillations and an increase in bunch size. On the other side, for positive ADTS coefficients a partial beam loss is observed when crossing the threshold. For values from zero up to $6000/\mathrm{m}$ more than 40 and up to 90 percent of the beam current is lost. For higher positive ADTS coefficients the loss goes down close to zero again. So, for ADTS coefficients up to $6000/\mathrm{m}$ there is a difference in the observed behaviour for a positive and negative sign of the ADTS coefficient. While at negative coefficients the instability is self-containing, for positive coefficients a partial beam loss is observed until the beam stabilizes again.
To investigate this difference in behaviour above the threshold the time domain signal of the center-of-mass oscillation and the bunch size was studied in measurement and simulation.
IV.3 Bunch Position and Size
Figure 8: Synchronized measurement of vertical spot profile and center-of-mass amplitude as function of time. The image shows the vertical light spot measured at different points in the sawtooth like dynamic. In red the envelope of the center-of-mass motion is displayed with a point for each spot profile measurement. The yellow dots indicate the spot size gained from a Gaussian fit. Measured at ADTS coefficient $b=-10000$ and current of 2.6 mA.
For a negative ADTS coefficient the dynamic above threshold shows clear, regular, sawtooth like bursts in bunch size and as amplitude modulation of the center-of-mass oscillations. For measurements this is visible in the BPM trace directly (Fig. 2) as well as in the synchronous measurement of bunch position and bunch size in Fig. 8. The contribution of the bunch size can be seen in the fact that the spot size goes down slower than the center-of-mass oscillations seen by the BPMs. Additionally, the calculated spot size of up to 1.7 mm (determined by Gaussian fit) in Fig. 8 compared to the maximal detected center-of-mass oscillation amplitude of 0.25 mm (= 0.5 mm peak-peak) indicates that the bunch size has a non-negligible contribution. Furthermore, the center-of-mass oscillation goes back to nearly zero for some hundred turns between the increases in oscillation amplitude. During this time the observed spot size slowly damps down indicating that this corresponds to the bunch size damping down, reaching minimal values in the order of 0.3 mm before the next peak. The same behaviour is present in the simulation shown in Fig. 9.
Figure 9: Simulated center-of-mass oscillation and bunch size at an ADTS coefficient $b=-15000$ and a current of 4.8 mA.
Figure 10: Top: Measured center-of-mass oscillation at 3.07 mA (left) and 2.79 mA (right) and an ADTS coefficient $b=13720$/m. The measurement at high current show a similar sawtooth pattern to measurements at negative ADTS coefficient (Fig. 2). The measurement at low current is more constant (noise level would be +-0.02 mm). Bottom: Simulated center-of-mass oscillation and bunch size at $b=15000$/m and bunch currents of 4.8 mA (left) and 4.6 mA (right).
This dynamic indicates a stabilizing mechanism which leads to a containment of the instability instead of a continuous growth until charge is lost. The observed behavior in the amplitude of the center-of-mass oscillations and the bunch size show that at some point a temporary stabilization occurs which leads to a damping down of the oscillation to below the noise limit of the measurement. The bunch size is also damped down during this stable period but it does not reach the expected stable bunch size before the instability is triggered again leading to a fast blow up of the bunch size and the onset of strong center-of-mass oscillations. A possible mechanism could be via an increased tune spread due to the blown up bunch size and the ADTS leading to an increased Landau damping effect. While the IBS could be the cause of the hysteresis observed in the Instability threshold, it is not a candidate for explaining the self-containing dynamics, as it operates in the wrong sense, ie. a blown-up beam has a lower threshold current and is therefore more unstable and does not contribute to a self-containing effect, where the threshold would need to increase to temporarily stabilize the unstable beam.
For positive ADTS, the dynamics above the threshold can only be observed in measurements at high ADTS coefficients where no instantaneous charge loss occurs. For higher bunch currents the dynamics in the center-of-mass oscillation and the bunch size have a similar sawtooth like pattern as observed for negative ADTS coefficients (see left side of Fig. 10). The pattern changes for lower bunch currents as shown on the right side in Fig. 10. Here, the center-of-mass oscillation amplitude is more constant and lightly modulated in the measurements, and nearly constant in the simulations.
Comparing the simulated maximal oscillation amplitudes and bunch sizes reached in, for example, Fig. 9 and the lower left plot of Fig. 10, shows a stronger blow up of the beam for the positive sign of the ADTS. The same difference is observed in measurements when comparing the maximal amplitude of the center-of-mass oscillation in Fig. 2 and the upper left plot in Fig. 10.
This indicates, that the level of “blow up” at which the instability is contained and finds some kind of equilibrium, pseudo-stable state is different for negative and positive ADTS. The asymmetry is visible clearly in Fig. 11, where for a bunch current slightly above threshold, the maximal bunch size and the maximal oscillation amplitude of the center-of-mass is given as a function of ADTS coefficient for simulations [^1] To be more robust against outliers the 95th percentile of the bunch size and the center-of-mass oscillation amplitude are taken.. The range in ADTS coefficient where partial current loss would occur probably depends on the combination of the center-of-mass oscillations and the total bunch size which, above a certain value, would lead to parts of the charge being “scraped” by the beam pipe. Independent of the exact value, it can be seen from Fig. 11 that the affected ADTS range would not be symmetric around zero but rather shifted to positive ADTS coefficients. In measurements, the same dependence of the maximal center-of-mass oscillation amplitude [^2] Again, the 95th percentile of center-of-mass oscillation is taken. as a function of the ADTS coefficient is observed (Fig. 12) for negative ADTS. While it is not measurable at the lower positive ADTS coefficients, due to the partial beam losses, the measured values at higher positive ADTS coefficients are higher than the corresponding values at negative ADTS showing the same asymmetry as in the simulations in Fig. 11. So to reach the same level of suppression of the instability, meaning low values in maximal bunch size and center-of-mass oscillations, a higher positive than negative ADTS coefficient would be needed. This asymmetry observed in both measurement and simulations supports the explanation for the asymmetry observed in the beam loss at the threshold (see Fig. 7).
Figure 11: Simulated maximal (95percentile) bunch size and center-of-mass oscillation amplitude and minimal (5percentile) action $\langle J\rangle$ as a function of the ADTS coefficient for currents slightly above threshold. The lines highlight the $1/x^{1/2}$ (bunch size and COM) respective $1/x$ ($\langle J\rangle$ dependency with the dotted line being the mirror of the solid line at negative ADTS.
Figure 12: Measurements of the maximal center-of-mass oscillation amplitude (95percentile) as a function of the ADTS coefficient. The Measurements were taken at bunch currents close to the threshold and with a chromaticity of 0.05. The lines highlight the $1/x^{1/2}$ dependency with the dotted line being the mirror of the solid line at negative ADTS. The errors show the standard deviation between multiple consecutive measurements per point.
Figure 13: Tune shift calculated from the maximal (95percentile) and the minimal (5percentile) action $\langle J\rangle$ as function of ADTS coefficient.
Figure 11 also shows the value of the average action of the particle ensemble $\langle J\rangle$ at the times of minimal bunch size and center-of-mass oscillations [^3] The 5th percentile is taken as value for the minimal action $\langle J\rangle$. as a function of ADTS coefficient, again with the asymmetry for the different signs of the ADTS coefficient visible. The minimal value $\langle J\rangle$ reaches can be connected to the point were the instability is no longer damped and the beam becomes unstable again, as will be described further in the next section (Sec. IV.4).
Both, the maximal bunch size and center-of-mass oscillation amplitudes as well as the minimal action $\langle J\rangle$ show a characteristic dependence on the ADTS coefficient. For $\langle J\rangle$ it follows a $1/x$ dependency and the bunch size as well as the center-of-mass oscillation amplitude has a $1/\sqrt{x}$ dependency. This already hints at a connection with the tune shift via Equ. 3. Going one step further and calculating the tune shift due to ADTS from the simulated values of the minimal action $\langle J\rangle$ for each ADTS coefficient, shows a constant but different level of tune shift for each sign of the ADTS (Fig. 13). While for negative ADTS the calculated tune shift of $\approx 0.00012$ is approximately 8% of the synchrotron tune, the shift for positive ADTS is with $\approx 0.00094$ already 65% of the synchrotron tune. This significant difference gives rise to the hypothesis, that for the positive and negative ADTS a different level of tune shift is required to contain the instability. The tune shift calculated from the maximal $\langle J\rangle$ shows a very similar behavior, where the small difference between maximal and minimal values on the positive side of the ADTS is explained by the fact that nearly no sawtooth is observed for positive ADTS coefficients and currents close to the threshold (compare Fig. 10). Overall, it can be concluded that the tune shift stays between these two levels, with the higher one being the point were the beam stabilises and starts to damp and the lower one indicates when the stabilizing effect stops and the beam goes unstable and blows up again. This is supported in the next section by theoretical calculations evaluating the instability threshold of the transverse mode coupling taking into account the Landau damping.
IV.4 Theoretical Calculations
Refer to caption Refer to caption
Figure 14: Stability diagrams for negative ADTS coefficient (left) and positive ADTS coefficient (right). The numbers that label the Landau contours indicate the magnitude of the tune shift at $\langle J\rangle\times 10^{3}$. The colored points are the eigenvalues of the scaled coupling matrix $\nu_{s}\mathbf{M}_{nl}^{mk}$ where the color represents the bunch current.
Figure 14 shows stability diagrams calculated from Eq. 6 for the cases of positive and negative ADTS. Contours are drawn by plotting the imaginary part of the inverse dispersion relation $I^{-1}_{m}$ against the real part. These contours map out a teardrop shape in complex frequency space pointing towards negative coherent tune shifts in the case of negative ADTS. Changing the sign of the ADTS to positive reflects the contour about the line of zero coherent frequency shift. Also shown on the figure are the eigenvalues of the coupling matrix without Landau damping where two modes are included: the azimuthal head tail modes $m=0$ and $m=-1$. For a head-tail mode to be stable in isolation, its complex coherent frequency shift would have to be within the Landau contour. The condition in the presence of mode coupling is slightly different, given by the zero determinant in Eq. 6, but the images are still illustrative none the less. It can be seen that, in order to influence the stability, a positive ADTS coefficient must be much larger in magnitude than a negative one, matching the observations in measurement and simulation shown in Fig. 11 and 12. This is intuitive as a negative ADTS coefficient means that the tune spread is towards negative tune shifts and the current-dependent tune shift of the $m=0$ mode is also negative for most broadband impedances.
One feature of storage rings used for fourth-generation synchrotron light sources, particularly those using low-RF frequencies (such as the 100 MHz of the 3 GeV ring at MAX IV), that is beneficial in this regard is the low incoherent synchrotron frequency. This means that the coupling frequency of the $m=0$ and $m=-1$ head-tail modes is not so far out of the spread in betatron tune of the electron bunches. Nevertheless, both in simulation and measurement at the 3 GeV ring at MAX IV, the magnitude and sign of the ADTS coefficient does not impact the threshold current of the TMCI. As discussed in Sec. IV.1, this is because the point of the mode-coupling is outside the tune spread of the bunch when it is stable. When the bunch goes unstable and increases in size, however, the tune spread increases until the Landau damping kicks in. Figure 15
Figure 15: Predicted equilibrium actions $\langle J\rangle$ and those determined from simulations in Mbtrack2 for an ADTS coefficient $b=1000\,\mathrm{tm^{-1}}$.
shows the saturation equilibrium average particle actions $\langle J\rangle$ in simulation and from the results of Eq. 6 for the two azimuthal head-tail modes and for a given ADTS coefficient $b$. Radiation damping is included in the simulations so they show the same sawtooth behavior seen in the measurements. The saturation action $\langle J\rangle$ is then taken as the minimum in this sawtooth pattern, corresponding to the point where the damping stops and the bunch blows up again, because during the rest of the sawtooth period, the beam is either being damped (and is therefore stable) or has some centroid motion, which has a negative effect on the Landau damping (this is a potential root cause of the sawtooth behavior). The agreement is not perfect, particularly in the value of the threshold current. This may change with the number of head-tail modes included in the theoretical prediction, although including more significantly complicates the numerical optimisation.
IV.5 Betatron Tune Shift with Current
(a) Simulation at negative ADTS
(b) Simulation at positive ADTS
(c) Measurement at negative ADTS
(d) Measurement at positive ADTS
Figure 16: Coherent motion spectrum showing the current dependent betatron tune shift below and above the instability threshold. Simulation (top): Fourier transform of the center-of-mass oscillation plotted as a function of the bunch current for negative (left) and positive (right) ADTS coefficient $b=15000$/m. Measurement(bottom): Fourier transform of the center-of-mass oscillation as a function of bunch current for an ADTS coefficient of $b=-10000$/m (left) and $b=13720$/m. During the measurement the instability was “switched” on and off (see text).
Observing the vertical betatron tune as function of current directly shows the expected current-dependent tune shift due to the transverse impedance from the zero-current tune of $\nu_{0}=0.275$ towards the -1 mode at the first synchrotron frequency side band ($\nu_{0}-0.00146$). In simulations of the coherent beam spectrum the threshold of the TMCI is clearly visible as the current at which the tune couples to the -1 mode (top row in Fig. 16). This is the same for both signs of the ADTS. The difference for negative and positive ADTS starts above the threshold where for negative ADTS the tune seems to continue its shift towards a lower tune with a similar slope as below the threshold (Fig. 16a). For the positive ADTS, the behaviour looks very different. While a slight shift in the opposite direction to higher tunes would not be unexpected, due to the positive sign of the ADTS, the tune jumps within a very small current range above the threshold from the -1 mode back to the 0 mode and then shows a continuous shift to higher tunes from there (Fig. 16b).
This drastic difference in behaviour can also be seen in measurements. Figures 16c and 16d show the measured tune spectra at different bunch currents and for negative and positive ADTS respectively. The measurements were conducted in such a way that the previously described hysteresis of the instability threshold (Sec. IV.1) was used to get comparative measurements for the tune of a stable and an unstable beam. To this end, the measurement was started at high bunch currents and the tune spectrum was recorded alternately for a beam stabilized by the BBB feedback system [^4] After initial stabilization the feedback is switched off during the measurement. and for an unstable beam, where the instability was triggered by a short excitation [^5] The excitation is switched off as well before the measurement is taken..
For the stable beam, the tune continues its current-dependent shift towards lower values. For both signs of the ADTS, it is clearly visible that the presence of the instability shifts the tune compared to the tune of the stable beam. For negative ADTS (Fig. 16c), the shift is small and towards slightly lower tune values. For positive ADTS (Fig. 16d), the tune is shifted back towards the zero-current tune (0 mode) and shows a very small current-dependent shift towards higher tune. Except for the difference in threshold and the threshold hysteresis observed in the measurements, the simulation and the measurements agree very well with respect to the tune shifts below and above threshold.
At higher bunch currents, additional features appear. In the measurement at negative ADTS, an upper and lower sideband shows up, moving with the tune as function of current. In the case of the positive ADTS the tune peak is broadened greatly and nearly spans from the -1 mode to the +1 mode. Comparing with the calculated tune shift of $\approx 0.001$ resulting from the maximal action $J$ simulated in case of positive ADTS (Fig. 13), shows that the jump of the coherent betatron tune by one synchrotron tune ($\nu_{s}=0.00146$) towards the 0 mode is only slightly bigger. Overall, from the measurement and tracking simulations it is not apparent whether this difference in the behaviour of the coherent tune above threshold is the cause or a consequence of the observed asymmetry in the level of beam “blow up” for negative versus positive ADTS.
V summary and conclusion
Landau damping has been investigated in the past as a possible source as mitigation mechanism of mode-coupling instabilities, also in connection with the amplitude-dependent tune shift as the source of the required tune spread.
While during standard operations the bunch current at the 3 GeV ring at the MAX IV Laboratory is below the TMCI threshold, an asymmetric dependence on the sign of the amplitude-dependent tune shift (ADTS) has been previously observed in dedicated experiments. Systematic studies were now conducted to investigate this observed asymmetry in dedicated single bunch experiments. It was observed, that for some ADTS coefficients the beam was lost when crossing the threshold while at others a saw-tooth shaped amplitude modulation was observed on the center-of-mass oscillation as well as on the bunch size leading to a self-contained instability. The presented simulations with the tracking tool mbtrack2 and the conducted measurements are in good agreement. Both show that the observed threshold current is independent of the ADTS coefficient and an observed hysteresis in the measured threshold can be attributed to intra-beam scattering effects. For the dynamics above the threshold, both measurements and simulation, show that for positive ADTS coefficients the maximal center-of-mass oscillation amplitude and bunch size that is reached before the instability stabilizes is systematically higher than for negative ADTS coefficients, indicating that this could be the cause of the observed partial beam current losses. The same asymmetry is also visible in the tune shift at the minimal $\langle J\rangle$ required for damping. The shift is constant and the value is only dependent on the sign of the ADTS, with the tune shift calculated for positive ADTS coefficients already being at $\approx 65\%$ of the synchrotron tune. Stability diagrams with Landau contours, calculated to include the amplitude-dependent tune shift, show as well that higher positive ADTS coefficients compared to negative ones are required for the instability to be Landau damped. Furthermore, simulations and measurements of the coherent tunes as a function of the bunch current, show a strong difference in the tunes development above threshold. For negative ADTS coefficients the tune is slightly shifting to lower values starting from the -1 mode at the threshold. In contrast, for positive ADTS coefficients, the coherent tune jumps back to the 0 mode and only then shows with increasing current a slight shift to higher tune values, as might be expected for positive ADTS coefficients.
Overall, it can be concluded that, for the presented measurements, the sign of the ADTS leads to an asymmetry in how strongly the vertical TMCI is self-containing and in which tune shift with current is observed above the threshold. As expected, a higher absolute value of the ADTS coefficient leads to a lower maximal center-of-mass oscillation and bunch size blow up. But when comparing signs, a higher positive ADTS coefficient is required for the same amount of suppression of the instability than for a negative ADTS.
Compared to previous studies on using Landau damping together with ADTS to mitigate the TMCI, two differences were found for the presented investigations for the parameters at the 3 GeV ring at MAX IV. It is clear that, the Landau damping comes in as a stabilization mechanism only after the beam begins to go unstable, as the threshold was shown to not be dependent on the ADTS coefficient, leading to the observed saw-tooth pattern. The second result and rather interesting finding is that the maximal level that the center-of-mass motion and bunch size reaches before being contained by Landau damping is higher for a positive than a negative sign of the ADTS coefficient. This could not only be shown in the measurement but also in tracking simulations and in theoretical stability considerations including Landau damping and the ADTS as well.
A key contribution to making this asymmetry visible, might by the rather low synchrotron frequency at the 3 GeV ring due to the low momentum compaction factor, typical in fourth-generation light-source storage rings combined with the low RF frequency of 100 MHz. Both of these aspects lead to a synchrotron frequency that starts to be within the betatron tune spread of the bunch when it is blown up by the instability. This would indicate that the presented findings might become more relevant in new fourth-generation light-source storage rings with even more extreme parameters than MAX IV.
Acknowledgements.
The computations were enabled by resources provided by LUNARC.